# Combined P-value Functions for Compatible Effect Estimation and Hypothesis Testing in Drug Regulation

### v1

@@ Abstract
The two-trials rule in drug regulation requires statistically significant results from two pivotal trials to demonstrate efficacy. However, it is unclear how the effect estimates from both trials should be combined to quantify the drug effect. Fixed-effect meta-analysis is commonly used but may yield confidence intervals that exclude the value of no effect even when the two-trials rule is not fulfilled. We systematically address this by recasting the two-trials rule and meta-analysis in a unified framework of combined p-value functions, where they are variants of Wilkinson’s and Stouffer’s combination methods, respectively. This allows us to obtain compatible combined p-values, effect estimates, and confidence intervals, which we derive in closed-form. Additionally, we provide new results for Edgington’s, Fisher’s, Pearson’s, and Tippett’s p-value combination methods. When both trials have the same true effect, all methods can consistently estimate it, although some show bias. When true effects differ, the two-trials rule and Pearson’s method are conservative (converging to the less extreme effect), Fisher’s and Tippett’s methods are anti-conservative (converging to the more extreme effect), and Edgington’s method and meta-analysis are balanced (converging to a weighted average). Notably, Edgington’s confidence intervals asymptotically always include individual trial effects, while meta-analytic confidence intervals shrink to a point at the weighted average effect. We conclude that all of these methods may be appropriate depending on the estimand of interest. We implement combined p-value function inference for two trials in the R package twotrials, allowing researchers to easily perform compatible hypothesis testing and parameter estimation.
Keywords: Confidence interval, estimand, median estimate, meta-analysis, two-trials rule
@@ 1 Introduction
The “two-trials rule” in drug regulation requires “at least two adequate and well-controlled studies, each convincing on its own” for the demonstration of drug efficacy and subsequent market approval [1, p.3]. This criterion reflects the need for “substantiation” and “replication” of scientific results [2, p.8], and is typically implemented by requiring the p-values from the two trials to be statistically significant at the conventional (one-sided) $\alpha=0.025$ level. However, this procedure alone does not provide a combined effect estimate nor a confidence interval (CI), and it has been suggested to pool the estimates with fixed-effect meta-analysis for this purpose [3, 4, 5]. Yet, the meta-analytic CI and point estimate are not always compatible with the two-trials rule. The meta-analytic CI may exclude the null value while the two-trials rule is not fulfilled, leading to discrepancies that are difficult to interpret and communicate.
The results from the two RESPIRE trials [6, 7, 8] in Table 1 illustrate this phenomenon. While the p-value for the null hypothesis of no effect from RESPIRE 1 is $p=0.004<0.025$, the p-value from RESPIRE 2 is $p=0.144>0.025$. Hence, the two-trials rule is not fulfilled at $\alpha=0.025$. At the same time, the 95% CI for the log rate ratio based on combining the trials’ log rate ratio effect estimates with fixed-effect meta-analysis ranges from $-0.58$ to $-0.08$ and thus excludes the value of 0.
Table 1: Results from the RESPIRE trials regarding the effect of ciprofloxacin after 14 days for the treatment of non-cystic fibrosis bronchiectasis [6, 7, 8].
@@
RESPIRE 1 $-0.49$ $[-0.85,-0.13]$ $0.004$
RESPIRE 2 $-0.18$ $[-0.53,~{}~{}~{}0.16]$ $0.144$
Meta-analysis $-0.33$ $[-0.58,-0.08]$ $0.004$
@@
A first attempt at resolving the apparent paradox could be to realize that the confidence level of the CI does not align with the level of the implicit test underlying the two-trials rule. Since the two-trials rule decision is based on two independent tests at level $\alpha=0.025$, the overall test is at level $\alpha^{2}=0.000625$, thus one could instead take a $(1-2\,\alpha^{2})\times 100\%=99.875\%$ meta-analytic CI [9, 10]. For the RESPIRE trials, this would lead to a meta-analytic $99.875\%$ CI from $-0.71$ to $0.05$ which includes the value of 0 and hence aligns with the two-trials rule decision. However, the level $\alpha=0.025$ is arbitrary and it would be desirable to have a CI that is compatible with the two-trials rule for any level, which is still not the case. For example, for $\alpha=0.05$, the two-trials rule is still not fulfilled, while the $(1-2\,\alpha^{2})\times 100\%=99.5\%$ meta-analytic CI from $-0.66$ to $-0.01$ excludes zero.
@@
Despite the widespread use of the two-trials rule in regulatory decision-making [11], it remains unclear how point and interval estimation should be reconciled with it. This paper aims to resolve this issue with a new approach. The key idea is to look at both the two-trials rule and meta-analysis from the perspective of p-value functions [12, 13, 14, 15, 16] and p-value combination methods [17, 18, 19, 20, 21]. The two-trials rule can be understood as a combined p-value function based on the squared maximum of two p-values [22] which is a special case of Wilkinson’s combination method [23], while meta-analysis corresponds to the combined p-value function based on Stouffer’s p-value combination method [24] with suitable weights. Both can be used to obtain combined p-values for the null hypothesis of no effect, CIs, and point estimates. These quantities are compatible in the sense that the (two-sided) p-value for a null value is less than $\alpha$ if and only if the null value is excluded by the $(1-\alpha)\times 100\%$ CI, and that the point estimate is included in the CI at any confidence level $(1-\alpha)\in(0,1)$. However, as we will show, the two methods implicitly target different estimands, which explains their different behaviors, and highlights the need to choose the method depending on the scientific question and corresponding estimand of interest. Moreover, the combined p-value function pespective suggests considering alternative p-value combination methods, for example, Edgington’s method based on the sum of p-values [25] or Fisher’s method based on the product of p-values [26].
@@
This paper is organized as follows: We begin by summarizing the general theory of combined p-value functions (Section 2), followed by investigating combined p-value functions based on the two-trials rule (Section 2.1), meta-analysis (Section 2.2), Tippett’s method (Section 3.1), Fisher’s and Pearson’s methods (Section 3.2), and Edgington’s method (Section 3.3) in more detail. For each, we derive corresponding point and interval estimates and investigate their properties. Results from two pairs of clinical trials are analyzed to illustrate the characteristics of the methods (Section 4). Extensions to more than two trials are discussed in Section 5. The paper ends with concluding discussions, limitations, and an outlook for future research (Section 6). Appendix A illustrates our R package twotrials for conducting p-value function inference, while Appendix B provides additional technical details.
@@ 2 Combined p-value functions
$p_{i}(\mu)=\begin{cases}1-\Phi(Z_{i})&\text{for}~{}H_{1i}\colon\theta_{i}>\mu~{}\text{(alternative = "greater")}\\ \Phi(Z_{i})&\text{for}~{}H_{1i}\colon\theta_{i}<\mu~{}\text{(alternative = "less")}\\ \end{cases}$ (1)
@@
cumulative distribution function of the standard normal distribution $\Phi(\cdot)$, null value $\mu$, and alternative hypothesis $H_{1i}$ chosen based on the orientation of the effect. For example, if a positive effect indicates treatment benefit, the alternative ”greater” would be chosen. We will not consider p-values with two-sided alternatives here, as the hypotheses tested in clinical trials usually have a well-defined direction. Moreover, combined p-value functions based on two-sided p-values can behave irregularly, e.g., they can be non-monotone so that the resulting confidence sets consist of empty or disjoint intervals, which is unintuitive and hard to communicate [27].
@@
$p(\mu)=g\left(p_{1}(\mu),p_{2}(\mu)\right)$
which combines the individual p-value functions $p_{1}(\mu)$ and $p_{2}(\mu)$ into a p-value function $p(\mu)$, which is a valid p-value function in the sense of having a uniform distribution for a particular $\mu$ if both $p_{1}(\mu)$ and $p_{2}(\mu)$ are also uniformly distributed for that $\mu$ [13, 27]. A two-sided $(1-\alpha)\times 100\%$ CI can then be obtained by determining the null values $\mu$ for which the p-value function is equal to $\alpha/2$ and $21-\alpha/$ 2. The so-called median estimate is given by the null value $\mu$ for which the p-value function equals 1/2 [28]. To obtain these quantities, it is useful to define a “combined estimation function”
@@
Figure 1: Illustration of a one-sided p-value function with alternative = ”greater” (left plot), corresponding two-sided p-value function (right plot), and corresponding 95% CI and median estimate.
In practice, it is informative to plot the p-value function for a range of null values $\mu$, see the left plot in Figure 1. For this purpose, it may also be converted to a two-sided p-value function using the transformation $2\min\{p(\mu),1-p(\mu)\}$, known as “centrality function” [13]. Such a two-sided p-value function then peaks at the median estimate, and it can be thresholded at $\alpha$ to conveniently read off the $(1-\alpha)\times 100\%$ CI [27], see the right plot in Figure 1.
When both trials have the same underlying true effect ($\theta_{1}=\theta_{2}=\theta$), sometimes called “one population assumption” or “homogeneity” [29, 30], a CI based on a combined p-value function has correct coverage and the median estimate is median unbiased for the true common effect $\theta$, i.e., the probability of the median estimate being greater than $\theta$ is equal to the probability of it being smaller than $\theta$ [see e.g., 13]. However, it is unclear how other operating characteristics (e.g., mean bias or CI width) behave for different combined p-value functions $g$, and how they behave when the true effects are not the same ($\theta_{1}\neq\theta_{2}$), known as “two populations assumption” or “heterogeneity” [29, 30]. In the following, we will investigate this in detail for the two-trials rule, meta-analysis, and four other types of combined p-value functions. As these investigations are somewhat technical, readers may choose to look only at the summary in Table LABEL:tab:summaries and then jump directly to the applications in Section 4.
@@ 2.1 The two-trials rule (maximum method)
The formulation using the squared maximum (2) may be preferable because $p_{\text{2TR}}(\mu)$ is a valid p-value, i.e., it has a uniform distribution if both $p_{1}(\mu)$ and $p_{2}(\mu)$ are also uniformly distributed for a particular $\mu$ [22]. The combined p-value function (2) is also a special case of Wilkinson’s p-value combination method based on the $r$th smallest out of $k$ p-values with $r=k=2$ [23]. This relationship can be used to generalize the two-trials rule to different settings while preserving type I error control at level $\alpha^{2}$, for example, settings with three rather than two trials [31]. We will discuss such extensions to more than two trials in Section 5 and focus first on effect estimation for two trials.
@@ 2.1.1 Effect estimation
Now assume that the standard errors of both trials are the same ($\sigma_{1}=\sigma_{2}=\sigma$) and the alternative is ”greater”. The median estimate is then
@@
where $\phi(\cdot)$ denotes the density function of the standard normal distribution, see Appendix B.1 for details. If the true effects from the two trials coincide ($\theta_{1}=\theta_{2}=\theta$) the first two terms of the expectation (6) reduce to the common effect $\theta$, whereas the last term reduces to $\sigma\times(z_{\sqrt{1/2}}-1/\sqrt{\pi})\approx\sigma\times-0.019$. Hence, the median estimate with (4) is negatively biased, yet the bias vanishes as the standard error decreases. In a similar way, one can show that the median estimate for the alternative ”less” is positively biased, so the median estimate from the two-trials rule exhibits a conservative bias in both cases.
Figure 2: Four hypothetical pairs of effect estimates and standard errors from two trials. The standard errors are assumed to be of the form $\sigma=\sqrt{2/n}$. The sample size $n$ in trial 1 is set to 5 (imprecise) or 500 (precise), and in trial 2 to twice the sample size of trial 1. The two-sided p-value functions of the individual trials (dashed lines), and the combined p-value functions (solid lines) based on the two-trials rule, fixed-effect meta-analysis, Tippett’s, Fisher’s, Pearson’s, and Edgington’s methods, and are shown along with the corresponding 95% CIs and median estimates (top). All p-values are based on the alternative ”greater” and then converted to two-sided p-values via the centrality function $2\min\{p,1-p\}$.
@@ 2.1.2 Asymptotics
see Appendix B.2 for details. Hence, the median estimate ($a=1/2$) and any CI limit ($a\neq 1/2$) approach $\min\{\hat{\theta}_{1},\hat{\theta}_{2}\}$ or $\max\{\hat{\theta}_{1},\hat{\theta}_{2}\}$, depending on the alternative hypothesis. This means that the CI shrinks to the more conservative of the two effects, while in case they coincide ($\theta_{1}=\theta_{2}=\theta$) it shrinks to the common effect $\theta$. Both scenarios are illustrated in Figure 2: In case of very small standard errors and different effect estimates (panel D), the CI based on the two-trials rule (black) is tightly concentrated around the smaller effect estimate, while for identical effect estimates (panel C), it is tightly concentrated around the common effect estimate.

### v2

@@ Abstract
[Abstract]The two-trials rule in drug regulation requires statistically significant results from two pivotal trials to demonstrate efficacy. However, it is unclear how the effect estimates from both trials should be combined to quantify the drug effect. Fixed-effect meta-analysis is commonly used but may yield confidence intervals that exclude the value of no effect even when the two-trials rule is not fulfilled. We systematically address this by recasting the two-trials rule and meta-analysis in a unified framework of combined p-value functions, where they are variants of Wilkinson’s and Stouffer’s combination methods, respectively. This allows us to obtain compatible combined p-values, effect estimates, and confidence intervals, which we derive in closed-form. Additionally, we provide new results for Edgington’s, Fisher’s, Pearson’s, and Tippett’s p-value combination methods. When both trials have the same true effect, all methods can consistently estimate it, although some show bias. When true effects differ, the two-trials rule and Pearson’s method are conservative (converging to the less extreme effect), Fisher’s and Tippett’s methods are anti-conservative (converging to the more extreme effect), and Edgington’s method and meta-analysis are balanced (converging to a weighted average). Notably, Edgington’s confidence intervals asymptotically always include the individual trial effects, while meta-analytic confidence intervals shrink to a point at the weighted average effect. We conclude that all of these methods may be appropriate depending on the estimand of interest. We implement combined p-value function inference for two trials in the R package twotrials, allowing researchers to easily perform compatible hypothesis testing and effect estimation.
keywords:
Confidence interval, estimand, median estimate, meta-analysis, two-trials rule
articletype: Research Article
@@ 1 Introduction
The “two-trials rule” in drug regulation requires “at least two adequate and well-controlled studies, each convincing on its own” for the demonstration of drug efficacy and subsequent market approval 1 p.3. This criterion reflects the need for “substantiation” and “replication” of scientific results 2 p.8, and is typically implemented by requiring the p-values from the two trials to be statistically significant at the conventional (one-sided) $\alpha=0.025$ level. However, this procedure alone does not provide a combined effect estimate nor a confidence interval (CI), and it has been suggested to pool the estimates with fixed-effect meta-analysis for this purpose 3, 4, 5. Yet, the meta-analytic CI and point estimate are not always compatible with the two-trials rule. The meta-analytic CI may exclude the null value while the two-trials rule is not fulfilled, leading to discrepancies that are difficult to interpret and communicate.
The results from the two RESPIRE trials 6, 7, 8 in Table 1 illustrate this phenomenon. While the p-value for the null hypothesis of no effect from RESPIRE 1 is $p=0.004<0.025$, the p-value from RESPIRE 2 is $p=0.144>0.025$. Hence, the two-trials rule is not fulfilled at $\alpha=0.025$. At the same time, the 95% CI for the log rate ratio based on combining the trials’ log rate ratio effect estimates with fixed-effect meta-analysis ranges from $-0.58$ to $-0.08$ and thus excludes the value of 0.
Table 1: Results from the RESPIRE trials regarding the effect of ciprofloxacin after 14 days for the treatment of non-cystic fibrosis bronchiectasis 6, 7, 8.
@@
RESPIRE 1 $-0.49$ $-0.85$ to $-0.13$ $0.004$
[…]
Meta-analysis $-0.33$ $-0.58$ to $-0.08$ $0.004$
@@
A first attempt at resolving the apparent paradox could be to realize that the confidence level of the CI does not align with the level of the implicit test underlying the two-trials rule. Since the two-trials rule decision is based on two independent tests at level $\alpha=0.025$, the overall test is at level $\alpha^{2}=0.000625$, thus one could instead take a $(1-2\,\alpha^{2})\times 100\%=99.875\%$ meta-analytic CI 9, 10. For the RESPIRE trials, this would lead to a meta-analytic $99.875\%$ CI from $-0.71$ to $0.05$ which includes the value of 0 and hence aligns with the two-trials rule decision. However, the level $\alpha=0.025$ is arbitrary and it would be desirable to have a CI that is compatible with the two-trials rule for any level, which is still not the case. For example, for $\alpha=0.05$, the two-trials rule is still not fulfilled, while the $(1-2\,\alpha^{2})\times 100\%=99.5\%$ meta-analytic CI from $-0.66$ to $-0.01$ excludes zero.
@@
Despite the widespread use of the two-trials rule in regulatory decision-making 11, it remains unclear how point and interval estimation should be reconciled with it. This paper aims to resolve this issue with a new approach. The key idea is to look at both the two-trials rule and meta-analysis from the perspective of p-value functions 12, 13, 14, 15, 16 and p-value combination methods 17, 18, 19, 20, 21. The two-trials rule can be understood as a combined p-value function based on the squared maximum of two p-values 22 which is a special case of Wilkinson’s combination method 23, while meta-analysis corresponds to the combined p-value function based on Stouffer’s p-value combination method 24 with suitable weights. Both can be used to obtain combined p-values for the null hypothesis of no effect, CIs, and point estimates. These quantities are compatible in the sense that the (two-sided) p-value for a null value is less than $\alpha$ if and only if the null value is excluded by the $(1-\alpha)\times 100\%$ CI, and that the point estimate is included in the CI at any confidence level $(1-\alpha)\in(0,1)$. However, as we will show, the two methods implicitly target different estimands, which explains their different behaviors, and highlights the need to choose the method depending on the scientific question and corresponding estimand of interest. Moreover, the combined p-value function pespective suggests considering alternative p-value combination methods, for example, Edgington’s method based on the sum of p-values 25 or Fisher’s method based on the product of p-values 26. All these p-value combination methods have been studied before in terms of hypothesis testing properties, such as admissibility or monotonicity 27, 17. In this paper, we take an alternative estimation perspective motivated by practical issues in drug regulation.
@@
This paper is organized as follows: We begin by summarizing the general theory of combined p-value functions (Section 2), followed by investigating combined p-value functions based on the two-trials rule (Section 2.1), meta-analysis (Section 2.2), Tippett’s method (Section 3.1), Fisher’s and Pearson’s methods (Section 3.2), and Edgington’s method (Section 3.3) in more detail. For each, we derive corresponding point and interval estimates and investigate their properties. Results from two pairs of clinical trials are analyzed to illustrate the characteristics of the methods (Section 4). Extensions to more than two trials are discussed in Section 5. The paper ends with concluding discussions, limitations, and an outlook for future research (Section 6). Appendix Combined P-value Functions for Compatible Effect Estimation and Hypothesis Testing in Drug Regulation illustrates our R package twotrials for conducting p-value function inference, while Appendix Combined P-value Functions for Compatible Effect Estimation and Hypothesis Testing in Drug Regulation provides additional technical details.
@@ 2 Combined p-value functions
$p_{i}(\mu)=\begin{cases}1-\Phi(Z_{i})&\text{for}\leavevmode\nobreak\ H_{1i}\colon\theta_{i}>\mu\leavevmode\nobreak\ \text{(alternative = "greater")}\\ \Phi(Z_{i})&\text{for}\leavevmode\nobreak\ H_{1i}\colon\theta_{i}<\mu\leavevmode\nobreak\ \text{(alternative = "less")}\\ \end{cases}$ (1)
@@
cumulative distribution function of the standard normal distribution $\Phi(\cdot)$, null value $\mu$, and alternative hypothesis $H_{1i}$ chosen based on the orientation of the effect. For example, if a positive effect indicates treatment benefit, the alternative "greater" would be chosen. We will not consider p-values with two-sided alternatives here, as the hypotheses tested in clinical trials usually have a well-defined direction. Moreover, combined p-value functions based on two-sided p-values can behave irregularly, e.g., they can be non-monotone so that the resulting confidence sets consist of empty or disjoint intervals, which is unintuitive and hard to communicate 28.
@@
$p(\mu)=g\left(p_{1}(\mu),p_{2}(\mu)\right),$
which combines the individual p-value functions $p_{1}(\mu)$ and $p_{2}(\mu)$ into a p-value function $p(\mu)$, which is a valid p-value function in the sense of having a uniform distribution for a particular $\mu$ if both $p_{1}(\mu)$ and $p_{2}(\mu)$ are also uniformly distributed for that $\mu$ 13, 28. A two-sided $(1-\alpha)\times 100\%$ CI can then be obtained by determining the null values $\mu$ for which the p-value function is equal to $\alpha/2$ and $21-\alpha/$ 2. The so-called median estimate is given by the null value $\mu$ for which the p-value function equals 1/2 29. To obtain these quantities, it is useful to define a “combined estimation function”
@@
Figure 1: Illustration of a one-sided p-value function with alternative = "greater" (left plot), corresponding two-sided p-value function (right plot), and corresponding 95% CI and median estimate.
In practice, it is informative to plot the p-value function for a range of null values $\mu$, see the left plot in Figure 1. For this purpose, it may also be converted to a two-sided p-value function using the transformation $2\min\{p(\mu),1-p(\mu)\}$, known as “centrality function” 13. Such a two-sided p-value function then peaks at the median estimate, and it can be thresholded at $\alpha$ to conveniently read off the $(1-\alpha)\times 100\%$ CI 28, see the right plot in Figure 1.
When both trials have the same underlying true effect ($\theta_{1}=\theta_{2}=\theta$), sometimes called “one population assumption” or “homogeneity” 30, 31, a CI based on a combined p-value function has correct coverage and the median estimate is median unbiased for the true common effect $\theta$, i.e., the probability of the median estimate being greater than $\theta$ is equal to the probability of it being smaller than $\theta$ (see e.g., Xie and Singh13). However, it is unclear how other operating characteristics (e.g., mean bias or CI width) behave for different combined p-value functions $g$, and how they behave when the true effects are not the same ($\theta_{1}\neq\theta_{2}$), known as “two populations assumption” or “heterogeneity” 30, 31. In the following, we will investigate this in detail for the two-trials rule, meta-analysis, and four other types of combined p-value functions. As these investigations are somewhat technical, readers may choose to look only at the summary in Table LABEL:tab:summaries and then jump directly to the applications in Section 4.
@@ 2.1 The two-trials rule (maximum method)
The formulation using the squared maximum (2) may be preferable because $p_{\text{2TR}}(\mu)$ is a valid p-value, i.e., it has a uniform distribution if both $p_{1}(\mu)$ and $p_{2}(\mu)$ are also uniformly distributed for a particular $\mu$ 22. The combined p-value function (2) is also a special case of Wilkinson’s p-value combination method based on the $r$th smallest out of $k$ p-values with $r=k=2$ 23. This relationship can be used to generalize the two-trials rule to different settings while preserving type I error control at level $\alpha^{2}$, for example, settings with three rather than two trials 32. We will discuss such extensions to more than two trials in Section 5 and focus first on effect estimation for two trials.
@@ 2.1.1 Effect estimation
Now assume that the standard errors of both trials are the same ($\sigma_{1}=\sigma_{2}=\sigma$) and the alternative is "greater". The median estimate is then
@@
where $\phi(\cdot)$ denotes the density function of the standard normal distribution, see Appendix Combined P-value Functions for Compatible Effect Estimation and Hypothesis Testing in Drug Regulation for details. If the true effects from the two trials coincide ($\theta_{1}=\theta_{2}=\theta$) the first two terms of the expectation (6) reduce to the common effect $\theta$, whereas the last term reduces to $\sigma\times(z_{\sqrt{1/2}}-1/\sqrt{\pi})\approx\sigma\times-0.019$. Hence, the median estimate with (4) is negatively biased, yet the bias vanishes as the standard error decreases. In a similar way, one can show that the median estimate for the alternative "less" is positively biased, so the median estimate from the two-trials rule exhibits a conservative bias in both cases.
Another interesting operating characteristic is the standard error of the median estimate. An intuitively desirable property is that the standard error of a combined estimate should not be larger than either of the trials’ standard errors. Assuming again that the true trial effects and standard errors coincide, the two-trials rule satisfies this, as the standard error takes the simple form
$\sigma_{\text{2TR}}=\sigma\sqrt{1-1/\pi}\approx\sigma\times 0.83,$ (7)
so is approximately $17\%$ smaller than the standard errors of each individual trial. However, this is no longer the case when the trial standard errors differ, where the standard error of the two-trial rule also depends on the true trial effects. See Appendix Combined P-value Functions for Compatible Effect Estimation and Hypothesis Testing in Drug Regulation for more details, also on standard errors based on other combined p-value functions.
Figure 2: Four hypothetical pairs of effect estimates and standard errors from two trials. The standard errors are assumed to be of the form $\sigma=\sqrt{2/n}$. The sample size $n$ in trial 1 is set to 5 (imprecise) or 500 (precise), and in trial 2 to twice the sample size of trial 1. The two-sided p-value functions of the individual trials (dashed lines), and the combined p-value functions (solid lines) based on the two-trials rule, fixed-effect meta-analysis, Tippett’s, Fisher’s, Pearson’s, and Edgington’s methods are shown along with the corresponding 95% CIs and median estimates (top). All p-values are based on the alternative "greater" and then converted to two-sided p-values via the centrality function $2\min\{p,1-p\}$.
@@ 2.1.2 Asymptotics
see Appendix Combined P-value Functions for Compatible Effect Estimation and Hypothesis Testing in Drug Regulation for details. Hence, the median estimate ($a=1/2$) and any CI limit ($a\neq 1/2$) approach $\min\{\hat{\theta}_{1},\hat{\theta}_{2}\}$ or $\max\{\hat{\theta}_{1},\hat{\theta}_{2}\}$, depending on the alternative hypothesis. This means that the CI shrinks to the more conservative of the two effects, while in case they coincide ($\theta_{1}=\theta_{2}=\theta$) it shrinks to the common effect $\theta$. Both scenarios are illustrated in Figure 2: In case of very small standard errors and different effect estimates (panel D), the CI based on the two-trials rule (black) is tightly concentrated around the smaller effect estimate, while for identical effect estimates (panel C), it is tightly concentrated around the common effect estimate.